# TW-CRL: Time-Weighted Contrastive Reward Learning for Efficient Inverse Reinforcement Learning

### v1

@@ 4.2.2 Contrastive Reward Learning
In TW-CRL, the reward function $r_{\phi}$ is modeled using a deep neural network and trained in a supervised manner. The training process involves creating a dataset from both successful and failed demonstrations. Building on the previously introduced Time-Weighted scheme, successful and failed demonstrations are converted into supervised learning labels for training the reward function. The key idea is to frame the estimation of $r_{\phi}(s_{t})$ as a supervised regression problem with contrastive labels, where successful demonstrations serve as “positive” examples and failed demonstrations as “negative” examples.
To create training labels, we use the Time-Weighted function $w(t)$ introduced in Equation 1 to compute the estimated reward $r_{\text{est}}$ for each state as follows:
@@
For a state $s_{t}$ in a successful demonstration $\tau^{+}$, we assign $w(t)$ as its estimated reward, meaning the state is more likely to help achieve the goal. For a state $s_{t}$ in a failed demonstration $\tau^{-}$, we assign $-w(t)$, meaning the state is more likely to cause failure.
To train the reward function, we introduce the Contrastive Reward Learning loss function, which minimizes the mean squared error (MSE) between the predicted rewards and the estimated reward of $s_{t}$:
@@
$=\frac{1}{N_{\text{goal}}}\sum_{\tau\in\mathcal{D}^{+}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})-r_{\text{est}}(s_{t})\right)^{2}$
$\quad+\frac{1}{N_{\text{trap}}}\sum_{\tau\in\mathcal{D}^{-}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})-r_{\text{est}}(s_{t})\right)^{2},$ (4)
@@
where $\mathcal{D}=\mathcal{D}^{+}\cup\mathcal{D}^{-}$ represents the collection of all states from both successful and failed demonstrations, and N is the total number of states. The terms $N_{\text{goal}}$ and $N_{\text{trap}}$ represent the number of states in successful and failed trajectories, respectively. By minimizing this Contrastive Reward Learning loss function, the reward function is trained to distinguish between states associated with successful task completion and those leading to failure.
Why This Approach Works We now explain how we initially developed this approach and why it successfully learns the desired reward function. Intuitively, we aim to train a reward function that computes the reward of a state as:
which is the intended objective of the trained reward function $r_{\phi}$. However, in practice, we assign labels as described in Equation 2, which can be interpreted as:
$r_{\text{est}}(s_{t})=\begin{cases}w(t)\propto P(s_{t}\in\mathcal{S}_{\text{goal}}),&\text{if }s_{t}\in\tau^{+}\\ -w(t)\propto-P(s_{t}\in\mathcal{S}_{\text{trap}}),&\text{if }s_{t}\in\tau^{-}\end{cases}$ (6)
where successful and failed demonstrations are labeled separately in the supervised learning model. This is because, in continuous environments, it is difficult to obtain the same state in both successful and failed demonstrations to estimate $P(s_{t}\in\mathcal{S}_{\text{goal}})$ and $P(s_{t}\in\mathcal{S}_{\text{trap}})$ for labeling. Furthermore, we can prove that Equation 2 is equivalent to Equation 5 during optimization of the reward function.
@@
Consider the derivative of Contrastive Reward Learning loss function (Equation 4):
@@
$r_{\phi}(s_{t})=\frac{P(s_{t}\in\mathcal{S}_{\text{goal}})-P(s_{t}\in\mathcal{S}_{\text{trap}})}{2}.$
@@
4.2.3 TW-CRL: Complete Algorithm Pipeline
TW-CRL follows an iterative process that consists of training a reward function, optimizing a policy, and collecting new data through agent-environment interaction. Similar to other IRL methods, TW-CRL requires only successful demonstrations at the beginning, while failed demonstrations are collected later during the agent’s interaction with the environment.
TW-CRL estimates the reward for each state, $r_{\text{est}}(s_{t})$, using the Time-Weighted function (Equation 1) and optimizes the reward function $r_{\phi}$ by minimizing the Contrastive Reward Learning loss (Equation 4). Once the reward function is trained, any suitable reinforcement learning algorithm can be used to update the policy, represented as $\text{POLICY\text{-}OPT}(r,\pi)$, to optimize the policy $\pi_{\theta}$. After the initial policy is trained, the agent interacts with the environment and generates new trajectories, which may include both successful and failed demonstrations. These new demonstrations are added to the dataset to further improve the reward function. After updating the reward function, the policy $\pi_{\theta}$ is retrained based on the updated rewards.
This iterative process allows the policy to make better use of the improved reward signal, helping the agent focus on high-reward states while avoiding potential trap states. An overview of the TW-CRL framework is presented in Figure 1, and the corresponding pseudocode is provided in Algorithm 1.
@@ 4.2.3 TW-CRL: Complete Algorithm Pipeline
  Input: Initial reward function $r_{\phi}$, policy $\pi_{\theta}$, policy updater $\text{POLICY\text{-}OPT}(r,\pi)$, expert demonstrations $\mathcal{D}_{\text{e}}$.
  // Initialize empty lists $R_{\phi},R_{\text{est}}$
  $R_{\phi}\leftarrow[\,]$, $R_{\text{est}}\leftarrow[\,]$
  $\mathcal{D}\leftarrow\mathcal{D}_{\text{e}}$
  while convergence is not achieved do
     // Update reward function $r_{\phi}$
     for each state $s_{t}\in\mathcal{D}$ do
        $w(t)=\frac{t}{T}\exp\left(k\frac{t}{T}\right)$
        if $s_{t}$ in successful trajectories  then
           $r_{\text{est}}(s_{t})=w(t)$
        else
           $r_{\text{est}}(s_{t})=-w(t)$
        end if
        Append $r_{\text{est}}(s_{t})$, $r_{\phi}(s_{t})$ to $R_{\text{est}}$, $R_{\phi}$
     end for
     $\mathcal{L}_{\text{CRL}}=\frac{1}{N}\sum_{i=1}^{|R_{\phi}[i]|}\left(R_{\phi}[i]-R_{\text{est}}[i]\right)^{2}$
     Update reward function $r_{\phi}$ with $\mathcal{L}_{\text{CRL}}$
     // Update policy $\pi_{\theta}$
     $\theta\leftarrow\text{POLICY\text{-}OPT}(r_{\phi},\pi_{\theta})$
     // Collect new successful and failed demos $\mathcal{D}^{+}$, $\mathcal{D}^{-}$
     $\mathcal{D}=\mathcal{D}\cup\mathcal{D}^{+}\cup\mathcal{D}^{-}$
  end while

### v2

@@ 4.2 Contrastive Reward Learning
In TW-CRL, the reward function $r_{\phi}$ is modeled using a deep neural network and trained in a supervised manner. The training process involves creating a dataset from both successful and failed demonstrations. Concretely, successful and failed demonstrations are converted into supervised learning labels using the Time-Weighted function introduced previously. The key idea is to frame the estimation of $r_{\phi}(s_{t})$ as a supervised regression problem with contrastive labels, where successful demonstrations serve as “positive” examples and failed demonstrations as “negative” examples.
Intuitively, we aim to train a reward function with labels $r_{\text{est}}(s_{t})$ of the following form for each state $s_{t}$:
However, in continuous environments, it is difficult to obtain the same state in both successful and failed demonstrations to estimate $P(s_{t}\in\mathcal{S}_{\text{goal}})$ and $P(s_{t}\in\mathcal{S}_{\text{trap}})$ for labeling. To work around this issue, we instead use the Time-Weighted function $w(t)$ introduced in Section 4.1 to compute the estimated reward $r_{\text{est}}$ for each state as follows:
@@
For a state $s_{t}$ in a successful demonstration $\tau^{+}$, we assign a positive value $w(t)$ as its estimated reward, meaning the state is more likely to help achieve the goal. For a state $s_{t}$ in a failed demonstration $\tau^{-}$, we assign a negative reward $-w(t)$, meaning the state is more likely to cause failure. We show in Appendix A.4 that training the reward function using labels shown in Equation 10 is similar to Equation 9. We then introduce the Contrastive Reward Learning loss function to train the reward function in a simple way:
@@
$=\frac{1}{N}(\sum_{\tau\in\mathcal{D}^{+}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})-r_{\text{est}}(s_{t})\right)^{2}+\sum_{\tau\in\mathcal{D}^{-}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})-r_{\text{est}}(s_{t})\right)^{2})$ (12)
$=\frac{1}{N}(\sum_{\tau\in\mathcal{D}^{+}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})-w(t)\right)^{2}+\sum_{\tau\in\mathcal{D}^{-}}\sum_{s_{t}\in\tau}\left(r_{\phi}(s_{t})+w(t)\right)^{2}),$ (13)
@@
where $\mathcal{D}=\mathcal{D}^{+}\cup\mathcal{D}^{-}$ represents the collection of all trajectories from both successful and failed demonstrations, and $N$ is the total number of states. This loss function minimizes the Mean Squared Error (MSE) between the predicted rewards and the estimated reward of $s_{t}$. By minimizing the proposed loss function, the reward function is able to effectively learn to distinguish between states associated with successful task completion and those leading to failure.
4.3 TW-CRL: Complete Algorithm Pipeline
An overview of the TW-CRL framework is presented in Figure 1, and the corresponding pseudocode is provided in Algorithm 1. TW-CRL follows an iterative process that consists of training a reward function, optimizing a policy, and collecting new data through agent-environment interaction. Similar to other IRL methods, TW-CRL requires only successful demonstrations at the beginning, while failed demonstrations are collected later during the agent’s interaction with the environment.
To begin with, TW-CRL estimates the reward for each state, $r_{\text{est}}(s_{t})$, using the Time-Weighted function (Equation 8) and optimizes the reward function $r_{\phi}$ by minimizing the Contrastive Reward Learning loss (Equation 13). Once the reward function is trained, any suitable reinforcement learning algorithm $\text{POLICY\text{-}OPT}(r,\pi)$ can be used to update the policy $\pi_{\theta}$. We use TD3[10] in our experiments. After the initial policy is trained, the agent interacts with the environment and generates new trajectories, which may include both successful and failed demonstrations. These new demonstrations are added to the dataset to further improve the reward function. After updating the reward function, the policy $\pi_{\theta}$ is retrained based on the updated rewards. This iterative process allows the policy to make better use of the improved reward signal, helping the agent focus on high-reward states while avoiding potential trap states.
4.4 Extending TW-CRL to other task types
In the discussions above, we assume that the notion of “goal states” and “trap states” exist in the task. In other words, we assume that the task TW-CRL is solving must be an episodic task with terminal states that clearly signals success or failure. In this section, we discuss how TW-CRL could be applied to tasks beyond this category.
Extending to episodic tasks without clear success / failure terminal states. For this kind of tasks, we split all trajectories into success and failure according to the total episodic return. Specifically, we apply a threshold $r_{\theta}$, which is a hyperparameter, to all the expert demonstrations as well as collected trajectories. All trajectories with episodic returns higher than or equal to $r_{\theta}$ are classified as successful trajectories, while others are classified as failed trajectories.
Extending to continuous tasks. For this kind of tasks, we follow common practice and set a horizon $T$. Each trajectory is then truncated at $T$ timesteps. We can then treat it in the same way as episodic tasks without clear success / failure terminal states as described above. Since the value of $T$ is often defined in the environment (as in all the environments we use introduced in Section 5), there’s no need to tune it as a hyperparameter.
5 Experiments
In this section, we conduct experiments to evaluate the performance of TW-CRL (Section 5.1), visualize and analyze the reward function in TrapMaze-v1 (Section 5.2), examine TW-CRL’s ability to generalize to various goal settings (Section 5.3), and perform ablation studies (Section 5.4). Details regarding hyperparameters, network architectures, additional training setups, and other information are provided in Section C.1, Section C.2, and Section C.4.
Benchmarks We use eight environments in total. This includes U-Maze and its two variants TrapMaze-v1 and TrapMaze-v2 which feature unknown trap states and randomized goal locations[8, 5], HumanoidStand, Ant[36],MountainCarCountinuous[37], PandaReach, and PandaPush[11, 28]. The environments are illustrated in Section C.1.
Baselines We evaluate TW-CRL against six baselines: four IRL methods (GAIL [17], AIRL [9], MaxEnt [46] and MERIT [22]), ICRL [18], which leverages suboptimal demonstrations, and SASR [23], which incorporates failed demonstrations.
5.1 Comparison Evaulation
Figure 2 and Table 1 report the average episodic returns and their variances across all tasks and baselines. We provide five successful demonstrations in the basic U-Maze environment, while TrapMaze-v1 and TrapMaze-v2 each receive 35 successful demonstrations, all ending in a randomly initialized goal located within a single designated grid area of the maze. During training, the same grid area is used for goal initialization, though the specific location within that grid varies each episode. In PandaReach and PandaPush, we offer 10 successful demonstrations.
(a) U-Maze
(b) TrapMaze-v1
(c) TrapMaze-v1
(d) MountainCarCont.
(e) HumanStand
(f) Ant
(g) PandaReach
(h) PandaPush
Figure 2: Training curves on benchmarks. The solid curves represent the mean, and the shaded regions indicate the standard deviation over five runs.
Table 1: Average final return across benchmarks. The best performance is marked in bold, and ± represents the standard deviation over five runs.
Environment U-Maze TrapMaze-v1 TrapMaze-v2 MountainCarContinuous
TW-CRL 273.05 ± 1.21 253.12 ± 9.46 201.27 ± 7.47 94.23 ± 0.07
SASR 192.72 ± 78.15 196.28 ± 32.89 63.68 ± 26.71 93.97 ± 0.09
MERIT 194.72 ± 19.01 51.20 ± 43.68 68.83 ± 20.59 91.70 ± 0.19
ICRL 173.73 ± 25.25 108.12 ± 36.52 72.60 ± 26.35 92.02 ± 0.10
GAIL 138.64 ± 20.25 159.06 ± 25.53 113.94 ± 50.26 91.80 ± 0.39
AIRL 208.21 ± 17.83 160.22 ± 11.25 48.95 ± 19.92 93.86 ± 0.53
MaxEnt 161.28 ± 41.38 129.70 ± 18.50 77.89 ± 30.10 91.71 ± 0.17
Environment HumanStand (×$\times 10^{3}$) Ant (×$\times 10^{2}$) PandaReach PandaPush
TW-CRL 144.69 ± 1.05 24.26 ± 0.12 -5.98 ± 0.55 -34.89 ± 6.58
SASR 128.22 ± 22.9 17.62 ± 0.87 -47.41 ± 0.51 -96.77 ± 2.34
MERIT 135.39 ± 3.37 17.34 ± 1.15 -11.35 ± 19.85 -76.09 ± 14.64
ICRL 57.89 ± 7.67 18.14 ± 0.40 -24.00 ± 12.83 -88.32 ± 3.42
GAIL 121.32 ± 9.35 15.22 ± 3.98 -46.53 ± 2.28 -99.00 ± 1.40
AIRL 134.83 ± 10.45 18.72 ± 1.43 -10.76 ± 5.93 -97.39 ± 3.00
MaxEnt 116.63 ± 20.82 -0.12 ± 0.01 -5.97 ± 2.28 -96.45 ± 2.54
(a) TW-CRL
(b) GAIL
(c) AIRL
(d) MaxEnt
(e) SASR
The results indicate that the performance of TW-CRL is better than or comparable to all baselines in terms of average episode return, with faster convergence and reduced variance. For example, in the U-Maze environment, the average final return in our algorithm demonstrates enhancements of $41.71\%$, $40.22\%$, $57.16\%$, $96.96\%$, $31.15\%$, and $69.33\%$ over SASR, MERIT, ICRL, GAIL, AIRL, and MaxEnt, respectively.
5.2 Analysis of Learned Rewards
In this section, we analyze the learned reward functions across different methods using the TrapMaze-v1 environment to illustrate why TW-CRL outperforms other approaches. The simplified illustration of the TrapMaze-v1 environment map is shown in Figure 4(a), where the red area represents the goal states, the blue area represents the trap states, and the gray blocks represent walls that the agent cannot get through. We provide a total of 35 successful demonstrations in which agents follow the main path to the goal, as illustrated in Figure 4(b). Figure 3 visualizes the reward functions learned by different methods throughout the training process. An exhaustive figure containing all baseline methods is shown in Appendix D.2.
The reward function visualizations reveal that GAIL, AIRL, and MaxEnt fail to recognize unseen trap states, and give high rewards only to states seen in successful demonstrations. As a result, these methods limit the agent to only imitating expert behavior without discovering better ways to reach the goal. On the other hand, SASR can detect unseen trap states by using failed demonstrations. It primarily assigns higher rewards to states that frequently appear in successful trajectories and lower rewards to those that frequently appear in failed trajectories. However, this approach still limits the agent’s ability to explore and find better solutions.
TW-CRL overcomes these limitations by enabling the agent to avoid trap states and efficiently discover paths to the goal. This is achieved by 1) utilizing both successful and failed trajectories in the Contrastive Reward Learning process and 2) utilizing temporal information in our Time-Weighted function to only place high emphasis on later states that may have heavier impacts on the outcome of the trajectory. As a result, TW-CRL not only identifies both goal and trap states more accurately than imitation-only methods, but also encourages broader exploration—allowing the agent to uncover more efficient routes instead of confining themselves to expert demonstrations. For example, the shortcut in Figure 4(c) and the alternative paths in Figure 4(d) were never shown in the expert demonstrations, but TW-CRL allows the agent to explore these paths.
(a) TrapMaze-v1 map
(b) An expert demo
(c) A shortcut path
(d) Alternative paths
Table 2: Average return under three goal reset settings. 1, 3, and any. #goal represent for the number of grids where the goal state might be in.
# goal state(s) TW-CRL SASR MERIT ICRL GAIL
1 253.12 ± 9.46 196.28 ± 32.89 51.20 ± 50.95 108.12 ± 36.52 159.06 ± 25.53
3 240.48 ± 4.44 150.84 ± 50.95 96.37 ± 11.94 39.59 ± 17.93 59.79 ± 22.87
Any 221.33 ± 11.58 122.78 ± 20.39 67.93 ± 23.7 90.12 ± 22.56 55.25 ± 21.67
5.3 Generalization Analysis
We then evaluate TW-CRL’s ability to generalize beyond the goal configurations provided in the original demonstrations in the TrapMaze-v1 environment, where the goal is randomly placed within one or several grid cells. Specifically, we provide demonstrations where the goal is initialized within a single grid cell, but during training and testing, the goal is allowed to be initialized within 1, 3, or any grid cells in the maze. Table 2 reports the average returns for each configuration. While all methods experience some drop in performance as the distribution of possible goal states broadens, TW-CRL remains consistently superior to the baselines, suggesting that its learned reward function can adapt successfully to unseen goal locations. Further results and details can be found in Section C.1 and Section D.1.
5.4 Ablation Study
Effect of the Time-Weighted function Figure 5(a) and Figure 5(b) (left) compare TW-CRL with the Time-Weighted function (blue) to a simpler approach using constant rewards of $+1$ for success states and $-1$ for failure states (purple). Results show that using time-weighted rewards accelerates convergence and achieves higher final returns in both TrapMaze-v1 and TrapMaze-v2. The advantage is even clearer in TrapMaze-v2, where the time-weighted approach significantly outperforms the non-time-weighted version and vanilla TD3. These findings suggest that time weighting is particularly beneficial in complex tasks.
(a) TrapMaze-v1
(b) TrapMaze-v2
6 Conclusion
In this paper, we propose TW-CRL, an inverse reinforcement learning approach that infers reward functions by leveraging temporal information from both successful and failed demonstrations. By incorporating the Time-Weighted function and Contrastive Reward Learning, TW-CRL learns an accurate and informative reward function, enabling the agent to efficiently locate goals while avoiding repeated failures, thereby enhancing learning efficiency and increasing task success rates. Additionally, TW-CRL enhances exploration beyond expert demonstrations by discovering alternative high-value paths or shortcuts that were not explicitly demonstrated. Empirical evaluations demonstrate that TW-CRL outperforms baseline methods, showcasing its effectiveness in enhancing agent performance and learning efficiency in complex tasks.
Appendix A Mathematical Derivations
A.1 Derivation of Equation 5
Consider a trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),...,(s_{T-1},a_{T-1})\}$ with horizon $T$. Recall from Section 3.2 that:
$\mathcal{S}_{\text{trap}}=\{s_{\text{trap}}\in{\mathcal{S}-\mathcal{S}_{\text{goal}}}\mid\sum\limits_{s^{\prime}\in\mathcal{S}_{\text{trap}}}P(s^{\prime}\mid s_{\text{trap}},a)=1,\forall a\in\mathcal{A}\}.$ (14)
where ${\mathcal{S}}$ is the state space, $\mathcal{S}_{\text{goal}}$ is the set of all goal states, $\mathcal{S}_{\text{trap}}\subseteq{\mathcal{S}}$ is the set of all trap states. We assume the following:
  1. 1.
    The initial state $s_{0}\notin\mathcal{S}_{\text{trap}}$.
  2. 2.
    For each timestep $t=1,2,...,T-1$, $P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{t-1}\notin\mathcal{S}_{\text{trap}})=k$. For now, we assume that $k\in\mathbb{R}$ is a constant value.
Given these assumptions, we can model the dynamics of the agent as an absorbing Markov chain, where the trap states are absorbing states and non-trap states are transient states. We can derive the probability of the agent being in a trap state at timestep $t$:
$P(s_{t}\in\mathcal{S}_{\text{trap}})$ $=1-P(s_{t}\notin\mathcal{S}_{\text{trap}})$ (15)
$=1-P(s_{t-1}\notin\mathcal{S}_{\text{trap}})P(s_{t}\notin\mathcal{S}_{\text{trap}}|s_{t-1}\notin\mathcal{S}_{\text{trap}})$ (16)
$=1-P(s_{0}\notin\mathcal{S}_{\text{trap}})\prod\limits_{i=1}^{t}P(s_{i}\notin\mathcal{S}_{\text{trap}}|s_{i-1}\notin\mathcal{S}_{\text{trap}})$ (17)
$=1-\prod\limits_{i=1}^{t}P(s_{i}\notin\mathcal{S}_{\text{trap}}|s_{i-1}\notin\mathcal{S}_{\text{trap}})$ (18)
$=1-\prod\limits_{i=1}^{t}[1-P(s_{i}\in\mathcal{S}_{\text{trap}}|s_{i-1}\notin\mathcal{S}_{\text{trap}})]$ (19)
$=1-(1-k)^{t}$ (20)
In a failed demonstration, recall from Section 3.2 that we assume the final state is always a trap state ($s_{T}\in\mathcal{S}_{\text{trap}}$). Therefore, given the prior knowledge that a demonstration is a failed demonstration, we can derive the conditional probability of a state $s_{t}$ of timestep $t$ being a trap state as follows:
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})$ $=\frac{P(s_{t}\in\mathcal{S}_{\text{trap}},s_{T}\in\mathcal{S}_{\text{trap}})}{P(s_{T}\in\mathcal{S}_{\text{trap}})}$ (21)
$=\frac{P(s_{t}\in\mathcal{S}_{\text{trap}})}{P(s_{T}\in\mathcal{S}_{\text{trap}})}$ (22)
$=\frac{1-(1-k)^{t}}{1-(1-k)^{T}}$ (23)
We can extend the conclusion above to cases where $k$ is decided by the current timestep $t$ and horizon $T$. Formally,
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})$ $=\frac{1-(1-f(t,T))^{t}}{1-(1-f(t,T))^{T}}$ (24)
A.2 Derivation of Equation 6 and (7)
We first show how Equation 6 is derived. By applying Equation 5 to Equation 4 we have:
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})$ $=\frac{1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}}{1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{T}}$ (25)
To simplify computation, notice that $1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{T}\approx 1$ when $T$ is large. Therefore,
$P(s_{t}\in\mathcal{S}_{\text{trap}}|s_{T}\in\mathcal{S}_{\text{trap}})$ $\approx 1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (26)
Similarly, to derive Equation 7, we apply the same process for $P(s_{t}\in\mathcal{S}_{\text{goal}}|s_{T}\in\mathcal{S}_{\text{goal}})$:
$P(s_{t}\in\mathcal{S}_{\text{goal}}|s_{T}\in\mathcal{S}_{\text{goal}})$ $=\frac{1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}}{1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{T}}$ (27)
$\approx 1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (28)
A.3 Plotting the Time-Weighted function
Recall that in our experiments, we use the following function as our Time-Weighted function:
$w(t)=1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$ (29)
A plot of the above function with different values of $\alpha$ is shown in Figure 6. We can observe that as the value of $\alpha$ increases, the curve becomes steeper and moves towards later timesteps, indicating that the Time-Weighted function $w(t)$ places a higher significance on later timesteps. In practice, the value of $\alpha$ can be tuned as a hyperparameter.
Figure 6: Illustration of $w(t)$ with T=10 and various $\alpha$ values.
A.4 Justification for Equation 10
@@
Consider the derivative of the Contrastive Reward Learning loss function (Equation 13):
@@
$r_{\phi}(s_{t})=\frac{P(s_{t}\in\mathcal{S}_{\text{goal}})-P(s_{t}\in\mathcal{S}_{\text{trap}})}{2}.$ (33)
@@
Appendix B Pseudocode
@@
1:Input: initial reward $r_{\phi}$, policy $\pi_{\theta}$, updater $\textsc{PolicyOpt}(r,\pi)$, expert demos $\mathcal{D}_{\mathrm{e}}$
2:$R_{\phi}\leftarrow[\,]$,  $R_{\text{est}}\leftarrow[\,]$triangleright Initialize empty lists
3:$\mathcal{D}\leftarrow\mathcal{D}_{\mathrm{e}}$
4:while convergence not achieved do
5:     for each state $s_{t}\in\mathcal{D}$ do $\triangleright$ Estimate reward label for each state
6:         $w(t)=1-(1-\frac{e^{\alpha t}-1}{e^{\alpha T}-1})^{t}$
7:         if $s_{t}$ is in successful trajectories then
8:              $r_{\text{est}}(s_{t})=\phantom{-}w(t)$
9:         else
10:              $r_{\text{est}}(s_{t})=-\,w(t)$
11:         end if
12:         $R_{\text{est}}\!\leftarrow R_{\text{est}}\cup\{r_{\text{est}}(s_{t})\}$,  $R_{\phi}\!\leftarrow R_{\phi}\cup\{r_{\phi}(s_{t})\}$triangleright Append $r_{\text{est}}(s_{t})$, $r_{\phi}(s_{t})$ to $R_{\text{est}}$, $R_{\phi}$
13:     end for
14:     $\mathcal{L}_{\text{CRL}}=\dfrac{1}{N}\sum_{i=1}^{N}\!\bigl{(}R_{\phi}[i]-R_{\text{est}}[i]\bigr{)}^{2}$triangleright Update reward function $r_{\phi}$ with $\mathcal{L}_{\text{CRL}}$
15:     Update $r_{\phi}$ by minimising $\mathcal{L}_{\text{CRL}}$
16:     $\theta\leftarrow\textsc{PolicyOpt}(r_{\phi},\pi_{\theta})$triangleright Update policy $\pi_{\theta}$
17:     $\mathcal{D}\leftarrow\mathcal{D}\cup\mathcal{D}^{+}\cup\mathcal{D}^{-}$triangleright Collect new successful and failed demos $\mathcal{D}^{+}$, $\mathcal{D}^{-}$
18:end while
Appendix C Experiment
C.1 Experiment Environments